OpenStudy (anonymous):

Does the series tan 4/n converge or diverge?

OpenStudy (anonymous):

diverge. \[\sum_{n=1}^{\infty} \tan(4/n) = \sum_{n=1}^{\infty} \sin(4/n)/\cos(4/n)\] but we know that 0<=1sin(x)<=1 and cos 0<=cos <=1, so the limit doesn't =0, and hence fails according to the divergence test.

OpenStudy (anonymous):

wait what, sin (0) is 0 and cos (0) = 1 so the limit of a_n is 0.

OpenStudy (anonymous):

yes maybe, but the point is its a series. its an infinite amount of additions, and sin and cos oscillation between -1 and 1 (soz, it wasn't meant to be zero above, rather -1)

OpenStudy (anonymous):

so, it can't coverage. it oscillates, so the limit doesn't even exist.

OpenStudy (anonymous):

Can I use the comparison test? because I don't think that answer is enough for my final.

OpenStudy (anonymous):

compare with what?

OpenStudy (anonymous):

a smaller series so that if that smaller series diverges, the larger one diverges as well.

OpenStudy (anonymous):

dont belive so

OpenStudy (anonymous):

wolfram alpha says the comparison test, I just don't know how to justify it.

OpenStudy (anonymous):

I mean it doesn't tell me the comparison.

OpenStudy (anonymous):

acutally i think im wrong

OpenStudy (anonymous):

I got it. The comparison is to 1/n. Just divide, use Lhopitals.

OpenStudy (anonymous):

it does indeed diverge

OpenStudy (anonymous):

or use the identity sin n/n as n->0 is 1.

OpenStudy (anonymous):

ye true

OpenStudy (anonymous):

My mistake was I kept trying to justify how 1/n was smaller than tan 1/n but all I needed to do was divide.

OpenStudy (anonymous):

thanks for your help anyway.

OpenStudy (anonymous):

I appreciate the attempt nonetheless.

OpenStudy (anonymous):

hah np, im learning with you man. got my test in 1 month, so thanks for the question

OpenStudy (anonymous):

nice, add me to your friends list. we can help each other. my test is on monday and my final is 2 weeks afterwards.

OpenStudy (anonymous):

k, cool, tell me your answers lol/

OpenStudy (anonymous):

anway, g2g cya

OpenStudy (anonymous):

cya